korryn18:

What is the slope of the line that passes through the points (1, 1) and (–1, –5)?

dude:

^ To get that answer use the slope formula between 2 points \(\frac{\Delta y}{\Delta x}\) or \(\frac{y_2-y_1}{x_2-x_1}\) \(\large \frac{-5-1}{-1-1}=\frac{-6}{-2}=3\)